# GiFT: Gibbs Fine-Tuning for Code Generation

### v1

@@ Abstract
Training Large Language Models (LLMs) with synthetic data is a prevalent practice in code generation. A key approach is self-training, where LLMs are iteratively trained on self-generated correct code snippets. In this case, the self-generated codes are drawn from a conditional distribution, conditioned on a specific seed description. However, the seed description is not the only valid representation that aligns with its intended meaning. With all valid descriptions and codes forming a joint space, codes drawn from the conditional distribution would lead to an underrepresentation of the full description-code space. As such, we propose Gibbs Fine-Tuning (GiFT), a novel self-training method inspired by Gibbs sampling. GiFT allows self-generated data to be drawn from the marginal distribution of the joint space, thereby mitigating the biases inherent in conditional sampling. We provide a theoretical analysis demonstrating the potential benefits of fine-tuning LLMs with code derived from the marginal distribution. Furthermore, we propose a perplexity-based code selection method to mitigate the imbalanced long-tail distribution of the self-generated codes. Empirical evaluation of two LLMs across four datasets demonstrates that GiFT achieves superior performance, particularly on more challenging benchmarks.
@@
Haochen Li    Wanjin Feng    Xin Zhou    Zhiqi Shen Nanyang Technological University, Singapore haochen003@ntu.edu.sg
@@ 1 Introduction
Figure 2: Overview of GiFT. For each description $d_{i}$ in the seed dataset, we first translate it between descriptions and codes iteratively to draw codes from the marginal distribution based on the intention of $d_{i}$. Then, we calculate the perplexity of each generated code and employ weighted random sampling to select codes with codes from the tail being more likely to be selected for fine-tuning. Finally, all selected codes are paired with $d_{i}$ for fine-tuning.
@@
We evaluate GiFT on DeepSeek-Coder-6.7B Guo et al. (2024) and CodeLlama-7B Roziere et al. (2023) over APPS+ (Introductory-level and Interview-level) Dou et al. (2024), MBPP+ Liu et al. (2024a), and CodeInsight Beau and Crabbé (2024) datasets. Experimental results demonstrate the superiority of drawing codes from the marginal distribution instead of the conditional distribution, and perplexity-guided data selection benefits self-training over iterations.
@@ 3 Gibbs Fine-Tuning
Since $Var(\mathbb{E}_{d_{ij}}[c|d_{ij}])\geq 0$, the variance of codes drawn from the marginal distribution is greater than or equal to the expected variance of codes drawn from the conditional distribution conditioned on a certain $d_{ij}$ (e.g. the seed description $d_{i}$). Here the variance of $c$ reflects the diversity of self-generated codes. More diverse codes are found to benefit LLM fine-tuning Yuan et al. (2023).
@@
Simply selecting all correct codes is detrimental. On the one hand, we are more likely to sample more codes for easy descriptions in the seed dataset and less for harder ones. Fine-tuning with all codes will bias LLMs towards easy descriptions, so we only select $K$ codes for each description. For descriptions with fewer than $K$ codes, we resample existing codes to ensure balance. On the other hand, there is data imbalance in the code set of each $d_{i}$. According to Ding et al. (2024), the marginal distribution within each code set is found to follow a long-tail distribution. Employing random sampling to select $K$ codes makes codes from the tail occupy only a small proportion of the training data since they are seldom generated, which bias $\mathcal{L}_{marg}$ towards the head. In iterative self-training, this bias will be exacerbated where the knowledge distribution of the LLM shifts to be more peaked.
@@
$ppl|(c|d,\mathcal{M})=\exp\Big{(}-{\frac{1}{l}}\sum_{t=1}^{l}\log P_{\mathcal{M}}(c_{t}|c_{<t},d)\Big{)}$
@@
$w_{ij}=\frac{e^{\mathrm{ppl}(c_{ij})/T}}{\sum_{j=1}^{n_{i}}e^{\mathrm{ppl}(c_{ij})/T}}$
@@ Appendix A Prompting Templates for Code-to-Text and Text-to-Code Translation
An example of a prompt for code generation is shown in Tab. 3. We construct a template for code completion where we provide a function head and a function docstring. The function docstring could be replaced with descriptions from previous Gibbs sampling results. We additionally provide one input-output pair for MBPP-sanitized and CodeInsight.
An example of a prompt for code summarization is shown in Tab. 4. We provide in-context examples to help LLM learn summarization. The in-context example is selected from a pool. We construct the pool by asking LLMs to summarize the same dataset without in-context examples and then filter out meaningless or too-long responses. The code following “###Code:” could be replaced with codes from previous Gibbs sampling results.
Appendix B Algorithm Workflow
Appendix C Experimental Setup
C.1 Dataset Statistics
Dataset APPS+ APPS+ MBPP CodeInsight
Introductory Interview
Train 1,998 3,736 170 1,547
Test 90 367 378 1,856
Table 2: Statistics of the dataset used in our experiment.
C.2 Implementation Details
Appendix D More Experimental Results
D.1 Impact of RFT data in GiFT
D.2 Extended Experiments of $T$
D.3 Impact of $K$ in Data Selection

### v2

@@ Abstract
Training Large Language Models (LLMs) with synthetic data is a prevalent practice in code generation. A key approach is self-training, where LLMs are iteratively trained on self-generated correct code snippets. In this case, the self-generated codes are drawn from a conditional distribution, conditioned on a specific seed description. However, the seed description is not the only valid representation that aligns with its intended meaning. With all valid descriptions and codes forming a joint space, codes drawn from the conditional distribution would lead to an underrepresentation of the full description-code space. As such, we propose Gibbs Fine-Tuning (GiFT), a novel self-training method inspired by Gibbs sampling. GiFT allows self-generated data to be drawn from the marginal distribution of the joint space, thereby mitigating the biases inherent in conditional sampling. We provide a theoretical analysis demonstrating the potential benefits of fine-tuning LLMs with code derived from the marginal distribution. Furthermore, we propose a perplexity-based code selection method to mitigate the imbalanced long-tail distribution of the self-generated codes. Empirical evaluation of two LLMs across four datasets demonstrates that GiFT achieves superior performance, particularly on more challenging benchmarks. Source code is available at https://github.com/Alex-HaochenLi/GiFT.
@@
Haochen Li    Wanjin Feng    Xin Zhouthanks:   Corresponding author    Zhiqi Shen Nanyang Technological University, Singapore {haochen003, xin.zhou, zqshen}@ntu.edu.sg   wanjin_feng@outlook.com
@@ 1 Introduction
Figure 2: Overview of GiFT. For each description $d_{i}$ in the seed dataset, we first translate it between descriptions and codes iteratively to draw codes from the marginal distribution based on the intention of $d_{i}$. Then, we calculate the perplexity of each generated code and employ weighted random sampling to select codes with codes from the tail being more likely to be selected for fine-tuning. Finally, all selected codes are paired with $d_{i}$ for fine-tuning. The example shown in this figure is taken from MBPP-sanitized/6.
@@
We evaluate GiFT on DeepSeek-Coder-6.7B Guo et al. (2024) and CodeLlama-7B Roziere et al. (2023) over APPS+ (Introductory-level and Interview-level) Dou et al. (2024), MBPP+ Liu et al. (2024a), and CodeInsight Beau and Crabbé (2024) datasets. Experimental results demonstrate the superiority of drawing codes from the marginal distribution instead of the conditional distribution (+1.2% on MBPP+, +2.3% on CodeInsight, +9.8% on APPS+ dataset), and perplexity-guided data selection benefits self-training over iterations.
@@ 3 Gibbs Fine-Tuning
Since $Var(\mathbb{E}_{d_{ij}}[c|d_{ij}])\geq 0$, the variance of codes drawn from the marginal distribution is greater than or equal to the expected variance of codes drawn from the conditional distribution conditioned on a certain $d_{ij}$ (e.g. the seed description $d_{i}$). What’s more, $Var(\mathbb{E}_{d_{ij}}[c|d_{ij}])$ could become even larger if an LLM is sensitive to input descriptions, which further widens the gap between $Var(c)$ and $\mathbb{E}_{d_{ij}}[Var(c|d_{ij})]$. Here the variance of $c$ reflects the diversity of self-generated codes. More diverse codes are found to benefit LLM fine-tuning Yuan et al. (2023).
@@
Simply selecting all correct codes is detrimental. On the one hand, we are more likely to sample more codes for easy descriptions in the seed dataset and less for harder ones. Fine-tuning with all codes will bias LLMs towards easy descriptions Singh et al. (2023), so we only select $K$ codes for each description. For descriptions with fewer than $K$ codes, we resample existing codes to ensure balance. On the other hand, there is data imbalance in the code set of each $d_{i}$. According to Ding et al. (2024), the marginal distribution within each code set is found to follow a long-tail distribution. Employing random sampling to select $K$ codes makes codes from the tail occupy only a small proportion of the training data since they are seldom generated, which bias $\mathcal{L}_{marg}$ towards the head. In iterative self-training, this bias will be exacerbated where the knowledge distribution of the LLM shifts to be more peaked.
@@
$\mathrm{ppl}(c|d,\mathcal{M})=\exp\Big{(}-{\frac{1}{l}}\sum_{t=1}^{l}\log P_{\mathcal{M}}(c_{t}|c_{<t},d)\Big{)}$
@@
$w_{ij}=\frac{\exp(\mathrm{ppl}(c_{ij})/T)}{\sum_{j=1}^{n_{i}}\exp(\mathrm{ppl}(c_{ij})/T)}$
@@ 5 Experiments
Figure 4: Pass@1 (%) of applying RFT+RD with descriptions from a single step of Gibbs sampling to Deepseek-Coder-6.7B on MBPP+ and CodeInsight.
RFT+RD with Descriptions from a Single-step GiFT
Given that the rewritten descriptions are often incorrect, we evaluate the performance of another stronger setting for RFT+RD. That is, we use the self-generated descriptions from a single step of Gibbs sampling as the rewritten descriptions in RFT+RD. The results of DeepSeek-Coder-6.7B on MBPP+ and CodeInsight are shown in Figure 4. The improvement of stronger RFT+RD over the vanilla RFT+RD verifies our claim that it is better to leverage the outstanding ability of LLMs to translate the code back to a description instead of rewriting the description to a new one. However, we could see that GiFT still outperforms the stronger RFT+RD. We think this is because all of the descriptions are from the same code. Instead, in GiFT, descriptions are from various codes in each Gibbs sampling step.
@@
Figure 9: Pass@1 (%) of applying the proposed data selection method to RFT and GiFT on APPS+ (Introductory) and APPS+ (Interview). DS is short for data selection.
Ablation study for Perplexity-guided Data Selection
The proposed perplexity-guided data selection method can be applied not only to GiFT but also to baseline methods. The results of applying the proposed data selection method to RFT and GiFT with $T=2$ on DeepSeek-Coder-6.7B are shown in Figure 9. We can see that the data selection method does not bring significant improvement to RFT compared with GiFT.
Impact of $K$ in Data Selection
@@ 7 Conclusion
Acknowledgement
We thank the anonymous reviewers for their helpful comments and suggestions. This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
Limitations
There are mainly three limitations in this work. First, GiFT is only evaluated on LLMs with a size of around 7B across Python datasets. However, as we demonstrated in the theoretical analysis, the loss bias from the conditional distribution is independent of model sizes and programming languages. Thus, we expect that GiFT is also effective in bigger or smaller LLMs and other programming languages. Second, GiFT relies heavily on test cases to filter out wrong self-generated codes. In this paper, we mainly evaluate GiFT on datasets that already provide test cases in the training set. We do not evaluate GiFT on the most recent high-quality datasets like OSS-Instruct Wei et al. (2023). A possible solution is to ask LLMs to generate test cases and codes at the same time, which is studied by recent works Chen et al. (2022, 2024); Liu et al. (2024c). Third, GiFT outperforms RFT by translating between natural language descriptions and codes, which introduces additional code summarization steps in the self-generation process. In our future work, we will focus on improving the sampling efficiency of GiFT by introducing a dynamic sampling strategy. For example, we do not generate a fixed number of codes and select at most one correct code for the next round. Instead, we dynamically adjust the number of generated codes based on the estimated accuracy to ensure that at least one correct code is expected. Thus, LLMs tend to generate fewer code candidates for easier descriptions in each round.
@@ Appendix A Prompting Templates for Code-to-Text and Text-to-Code Translation
An example of a prompt for code generation is shown in Table 2. We construct a template for code completion where we provide a function head and a function docstring. The function docstring could be replaced with descriptions from previous Gibbs sampling results. We additionally provide one input-output pair for MBPP-sanitized and CodeInsight.
An example of a prompt for code summarization is shown in Table 3. We provide in-context examples to help LLM learn summarization. The in-context example is selected from a pool. We construct the pool by asking LLMs to summarize the same dataset without in-context examples and then filter out meaningless or too-long responses. The code following “###Code:” could be replaced with codes from previous Gibbs sampling results.
@@
Appendix B Algorithm Workflow
Appendix C Experimental Setup
C.1 Dataset Statistics
Dataset APPS+ APPS+ MBPP CodeInsight
Introductory Interview
Train 1,998 3,736 170 1,547
Test 90 367 378 1,856
Table 5: Statistics of the dataset used in our experiment.
C.2 Implementation Details
Appendix D More Experimental Results
D.1 Impact of RFT data in GiFT
D.2 Extended Experiments of $T$
Algorithm 1 Workflow of GiFT in each iteration
Input: A seed dataset $\mathcal{D}=\{d_{i}\}_{i=1}^{N}$, an LLM $\mathcal{M}$.
Parameter: Gibbs sampling iterations $n$, selection threshold $K$, temperature $T$.
Output: An updated LLM $\mathcal{M}^{*}$ for next GiFT iteration.
1:$\mathcal{C}\leftarrow\emptyset$.
2:for each description $d_{i}\in\mathcal{D}$ do $\triangleright$ Gibbs Sampling
3:     $c_{i1}\leftarrow\mathcal{M}(d_{i})$
4:     for $k\leftarrow 1$ to $n$ do
5:         $d_{ik}\leftarrow\mathcal{M}(c_{ik})$triangleright Summarize code into description
6:         $c_{ik+1}\leftarrow\mathcal{M}(d_{ik})$triangleright Generate code from description
7:     end for
8:     $\mathcal{C}_{i}\leftarrow\{c_{i1},c_{i2},\dots,c_{in}\}$
9:end for
10:$\mathcal{C}^{*}\leftarrow\emptyset$
11:for each description $d_{i}\in\mathcal{D}$ do $\triangleright$ Perplexity-Guided Selection
12:     for each code $c_{ij}\in\mathcal{C}_{i}$ do
13:         $ppl|(c_{ij})\leftarrow\exp\Big{(}-\frac{1}{|c_{ij}|}\sum_{t=1}^{|c_{ij}|}\log P_{\mathcal{M}}(c_{t}|c_{<t},d_{i})\Big{)}$triangleright Compute perplexity
14:     end for
15:     for each code $c_{ij}\in\mathcal{C}_{i}$ do
16:         $w_{ij}\leftarrow\frac{e^{\text{ppl}(c_{ij})/T}}{\sum_{j=1}^{n_{i}}e^{\text{ppl}(c_{ij})/T}}$triangleright Compute weight
17:     end for
[…]
19:     $\mathcal{C}_{i}^{*}\leftarrow\mathcal{C}_{i}^{*}\cup\{c_{ij}\}^{K}$
20:end for
21:$\mathcal{D}^{*}\leftarrow\{(d_{i},c)|d_{i}\in\mathcal{D},c\in\mathcal{C}_{i}^{*}\}$triangleright Construct dataset for SFT
22:$\mathcal{M}^{*}\leftarrow SFT(\mathcal{M},\mathcal{D}^{*})$triangleright Supervised Fine-Tuning
23:return $\mathcal{M}^{*}$triangleright Return fine-tuned model